# Mean-field games with rough common noise I: the linear-quadratic case

### v1

@@ 4.2 Connecting the Volterra and rough frameworks
Turning to the choice of the Volterra kernel, (S1-R) along with 2.9 guarantee that, for any $(m,m^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$, the deterministic controlled rough paths $\big(U^{\boldsymbol{\eta}},(U^{\boldsymbol{\eta}})^{\prime}\big)\in\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d\times d}\big)$ and $\big(M^{\boldsymbol{\eta},(m,m^{\prime})},(M^{\boldsymbol{\eta},(m,m^{\prime})})^{\prime}\big)\in\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ respectively solve the linear RDEs
@@
The matrix $U^{\boldsymbol{\eta}}_{t\leftarrow s}$ is invertible for every $t\in[0,T]$ (this follows e.g. by applying A.3). Its inverse $(U^{\boldsymbol{\eta}}_{t\leftarrow s})^{-1}=:U^{\boldsymbol{\eta}}_{s\leftarrow t}$, where $\big(U^{\boldsymbol{\eta}}_{s\leftarrow t},(U^{\boldsymbol{\eta}}_{s\leftarrow t})^{\prime}\big)\in\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d\times d}\big)$ is the solution of
@@
We shall need the following instance of Duhamel’s principle. From A.3, applied in ${U}^{\boldsymbol{\eta}}_{0\leftarrow t}M^{\boldsymbol{\eta},(m,m^{\prime})}_{t}$, we get that
@@
Now, let us denote with ${{X}}^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}$ the unique solution of (1), see 2.9, and with $\widetilde{{X}}^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}$ the unique solution, see 3.2, of equation (6) where $M(m)_{t}$ is replaced by $M^{\boldsymbol{\eta},(m,m^{\prime})}_{t}$ and
@@
Direct application of the rough product formula, A.3, to the product ${U}^{\boldsymbol{\eta}}_{0\leftarrow t}{X}^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}$ with ${U}^{\boldsymbol{\eta}}_{0\leftarrow t}$ satisfying (32).
@@
$\bigg(\mathcal{K}^{\boldsymbol{\eta}}(0,\cdot)C^{1}m\;,\;\mathcal{K}^{\boldsymbol{\eta}}(0,\cdot)^{\prime}C^{1}m+\mathcal{K}^{\boldsymbol{\eta}}(0,\cdot)(C^{1})^{\prime}m+\mathcal{K}^{\boldsymbol{\eta}}(0,\cdot)C^{1}m^{\prime}\bigg)\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{p\times d}\big).$ (34)
@@
Hence, for fixed $(m,m^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ and $\xi\in\bigcap_{k=2}^{\infty}\mathbb{L}^{k}(\mathcal{F}_{0},\mathbb{R}^{d})$, under (S1-R), (S2) and (S3), from the Volterra formulation where $M(m)$ is replaced from $M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)$, the stochastic optimal control problem has a unique solution. Furthermore, by 3.9 and 4.5, for the optimal state and optimal control $\big(X^{(m,m^{\prime}),\xi},\alpha^{*,(m,m^{\prime}),\xi}\big)$, of the rough stochastic optimal control problem we get the representations
@@
Indeed, by integrability of $\xi$ and the fact that the transformed problem (recall (19)) is solved through a (deterministic) decoupling field, standard SDE estimates, see e.g. [54, Theorem 3.4.3], imply
@@
We close this section with the analogue of 3.9.
@@
First we are going to establish existence of such a solution. From the discussion above, and specifically 3.9, there exist unique processes $X^{(m,m^{\prime}),\xi},Y^{(m,m^{\prime}),\xi},$ $(Z^{1})^{(m,m^{\prime}),\xi}$, with the appropriate integrabilities for $X^{(m,m^{\prime}),\xi},Y^{(m,m^{\prime}),\xi},$ (see 4.7) that solve (36). We are going to show that $X^{(m,m^{\prime}),\xi},Y^{(m,m^{\prime}),\xi}$ are also part of a solution for (37) along with an appropriate $Z^{(m,m^{\prime}),\xi}$.
@@
From 2.9 both equations admit unique solutions and an application of 4.5 yields $\mathbb{P}\big[X^{(m,m^{\prime}),\xi}=\widetilde{X}^{(m,m^{\prime}),\xi}\big]=1$.
@@
For the rest of the proof we are going to denote $P^{\boldsymbol{\eta}},\Pi^{\boldsymbol{\eta},(m,m^{\prime})}$ by $P,\Pi$ for ease of notation. We will apply A.3 several times in order to see that $Y^{(m,m^{\prime}),\xi}$ has the desired dynamics. Indeed, we have
@@
solve of (37) with the properties described in the statement of the theorem. By repeating the procedure, from the other direction, i.e. forward component first and then backward, with the help of A.3, we get that $\big(X^{(m,m^{\prime}),\xi},Y^{(m,m^{\prime}),\xi},\mathcal{K}^{\boldsymbol{\eta}}(\cdot,\cdot)^{\top}Z^{(m,m^{\prime}),\xi}\big)$ is a solution of (17). Notice that we are allowed to do that because, by definition, the solution space imposes sufficient integrability properties. In particular, $Z^{(m,m^{\prime}),\xi}$ is sufficiently integrable in order to apply A.3. Hence, from the uniqueness part of 3.9, uniqueness of (37) follows. ∎
@@ 4.3 Equivalence of rough fixed point step with FBRDE
From 4.5, with $\Sigma=0$ and $\xi$ deterministic (see the proof of 4.8), (39) is equivalent to the forward-backward RDE (FBRDE)
@@
Exactly the same as the proof of 4.2, with the usage of 4.5, or 4.8. ∎
@@ 4.4 Solving the FBRDE
From 4.10, the rough LQ MFG problem has a (unique) solution, provided that (43) has a (unique) solution.
@@
    We assume that $Q_{t}+\bar{Q}_{t}-\bar{Q}_{t}S_{t}\succeq 0$ and $Q+\bar{Q}-\bar{Q}S\succeq 0$. Additionally, let $(e,e^{\prime})\in\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d\times d}\big)$ be the unique solution to the linear RDE [^7] Uniqueness is standard, e.g. by straightforward extension of [31, Sec 8.9], or as special case of [12, Thm 3.11], or [42, Thm 3.5].
@@
Now, by A.2 we get a unique global solution for the symmetric rough Riccati
@@
Then, by the rough product formula, A.3, we get that $\bar{p}_{t}$ satisfies (45). So, $\bar{v}(t,x)$ is well defined [^8] Uniqueness of (45) comes by iterating local uniqueness.. Then, using the ansatz we decouple the first equation of (43) and write it as
@@
where $\bar{p}_{t}$ satisfies equation (45) and $\bar{x}_{t}$ satisfies (48). This can be proven immediately by using the rough product formula A.3 as above. Indeed, we have
@@
By virtue of A.3, and by repeating the proof of 4.8 (without worrying about integrability properties as we are in a deterministic setting), we see that (43) is equivalent to
@@ 5 Stability of the rough LQ MFG
$\rho_{\gamma}(\boldsymbol{\eta}_{1})+\rho_{\gamma}(\boldsymbol{\eta}_{2})\leq M$
then, up to a nonnegative multiplicative constant that depends on $M,T,\gamma,\beta,$
@@
$\rho_{\gamma}(\boldsymbol{\eta}_{1})+\rho_{\gamma}(\boldsymbol{\eta}_{2})\leq M$
then, up to a nonnegative multiplicative constant that depends on $M,T,\gamma,\beta,\xi,$ we have
@@
By classical stability estimates for rough integrals (see e.g. [31, Chapter 4]), the last term is bounded from above by
@@
almost surely for all $r\in[0,t].$ Next, notice that $P^{\boldsymbol{\eta}_{i}},\Pi^{\boldsymbol{\eta}_{i},(m_{i},m_{i}^{\prime})}$ are (differentiable-in-time) solutions of the deterministic system of backward matrix equations (27) with coefficients $\widehat{A}^{i},\widehat{B}^{i},\widehat{Q}^{i},\widehat{\bar{Q}}^{i},\mathcal{K}^{\boldsymbol{\eta}_{i}}$ depending on $\boldsymbol{\eta}_{i},i=1,2.$ Thus, up to multiplicative constants that depend polynomially on the seminorms $\rho_{\gamma}(\boldsymbol{\eta}_{i})$ we have the almost-sure estimate
@@ 6 Randomization
$(\mathbb{W}^{\perp}_{s,t})^{\mathrm{Strato}}:=\int_{s}^{t}\delta W^{\perp}_{s,r}\circ\mathrm{d}W^{\perp}_{r}$
It is well known (see [31, Chapter 3]) that for all $\omega^{\prime\prime}$ outside a $\mathbb{P}^{\prime\prime}$-null set $N^{\prime\prime}$ of $\Omega^{\prime\prime}$, the Brownian rough path
@@
Just to be clear, for $\omega^{\prime\prime}\in N^{\prime\prime}$ we set $\textbf{W}^{\perp}(\omega^{\prime\prime}):=(0,0)$, and the same convention will hold for every new process defined below whenever we randomize with respect to $\textbf{W}^{\perp}$.
@@
    There exist $A^{1},C^{1}:\mathscr{C}^{0,\gamma}_{g}\rightarrow C^{\beta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)$ measurable with respect to $\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)$ and $(A^{1})^{\prime},(C^{1})^{\prime}:\mathscr{C}^{0,\gamma}_{g}\rightarrow C^{\beta^{\prime}}\big([0,T];L\big(\mathbb{R}^{p};L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)\big)$ measurable with respect to $\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)$, such that $(A^{1}(\boldsymbol{\eta}),(A^{1})^{\prime}(\boldsymbol{\eta})),(C^{1}(\boldsymbol{\eta}),(C^{1})^{\prime}(\boldsymbol{\eta}))\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)\hskip 2.84544pt\text{for every}\hskip 2.84544pt\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}$, and
@@
    An example when assumption (S6-R) is satisfied is given in the case that $A^{1},C^{1}$ take values in $C^{\beta+\beta^{\prime}}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)$ and are constant with respect to $\mathscr{C}^{0,\gamma}_{g}$. Then, $(A^{1},0),$ $(C^{1},0)\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)$ for every $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$. For more examples one can apply the results of [31, Chapter 7].
Now, let $\xi\in\mathbb{R}^{d}$, $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$, $\alpha\in\mathcal{A}^{2}$ and $(m,m^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$. Then, under S6-R, from 2.9 and [42, Proposition 3.10], we get that there exists a unique prog. measurable with respect to $\big(\mathcal{F}_{t}\otimes\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)\big)_{t\in[0,T]}$ process $X^{\boldsymbol{\eta}}_{t}(\omega^{\prime}):=X^{\boldsymbol{\eta},\alpha(\boldsymbol{\eta}),(m(\boldsymbol{\eta}),m^{\prime}(\boldsymbol{\eta})),\xi}_{t}(\omega^{\prime})$ which solves
@@
For every $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}$, under (S6-R), there exist unique $(m^{*}(\boldsymbol{\eta}),(m^{*})^{\prime}(\boldsymbol{\eta}))\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ and $\alpha^{*}(\boldsymbol{\eta})\in\mathcal{A}$ such that
@@
From Lemma 5.2, we have that $m^{*}(\boldsymbol{\eta}),(m^{*})^{\prime}(\boldsymbol{\eta})$ depend continuously on $\boldsymbol{\eta}$, thus the decoupling coefficients $\mathcal{K}^{\boldsymbol{\eta}}(\cdot,\cdot),M^{\boldsymbol{\eta}}\big((m^{*}(\boldsymbol{\eta}),(m^{*})^{\prime}(\boldsymbol{\eta}))\big)_{\cdot},P^{\boldsymbol{\eta}}_{\cdot},\Pi^{\boldsymbol{\eta},(m^{*}({\boldsymbol{\eta}}),(m^{*})^{\prime}(\boldsymbol{\eta}))}_{\cdot}$ also depend continuously on $\boldsymbol{\eta}$ (and are deterministic). Moreover, we remind that $\mathscr{C}^{0,\gamma}_{g}$ is Polish, hence Carathéodory’s Lemma can be applied, see for instance [3, 4.51 Lemma]. So, $(m^{*},(m^{*})^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$, $X^{*,\boldsymbol{\eta}}$ is prog. measurable with respect to $\big(\mathcal{F}_{t}\otimes\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)\big)_{t\in[0,T]}$ and $\alpha^{*}\in\mathcal{A}^{2}$, hence we have proved the following:
@@
From the above [^12] See also [35, Theorem 3.6]. we get that $\overline{\alpha}\in\overline{\mathcal{A}},\hskip 2.84544pt\overline{m}\hskip 2.84544pt\text{is}\hskip 2.84544pt\mathcal{F}^{\perp}_{T}-\text{measurable},\hskip 2.84544pt\overline{m}^{\prime}\hskip 2.84544pt\text{is}\hskip 2.84544pt\mathcal{F}^{\perp}_{T}-\text{measurable}\hskip 2.84544pt,(\overline{m},\overline{m}^{\prime})\in\overline{\mathscr{M}}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})\hskip 2.84544pt\text{and}\hskip 2.84544pt\overline{X}\hskip 2.84544pt\text{is}\hskip 2.84544pt(\mathcal{F}_{t}\vee\mathcal{F}^{\perp}_{T})_{t\in\mathbb{R}_{+}}-\text{prog. measurable},$ where
@@
Moreover, by [34, Theorem 3.6], we get that for every $t\in[0,T]$ it holds
@@
so, by [28, Proposition 7.2] or [42, Proposition 5.2], we get for every $i\in\mathbb{N}$
@@
Before we state our final result, for $(m,m^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$, we define the following value functions.
@@
Because $\mathcal{A}^{1}\subseteq\mathcal{A}^{2}$, and by the first part of the proof of [28, Theorem 7.4], we have
@@
Let $(m^{*},(m^{*})^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$ , $\alpha^{*}\in\mathcal{A}^{2}$ be the objects described in 6.5. Then, by (63) and (64), we immediately get that
@@
Thus we arrive at the following result.

### v2

@@ 4.2 Connecting the Volterra and rough frameworks
Turning to the choice of the Volterra kernel, (S1-R) along with Theorem 2.9 guarantee that, for any $(m,m^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$, the deterministic controlled rough paths $\big(U^{\boldsymbol{\eta}},(U^{\boldsymbol{\eta}})^{\prime}\big)\in\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d\times d}\big)$ and $\big(M^{\boldsymbol{\eta},(m,m^{\prime})},(M^{\boldsymbol{\eta},(m,m^{\prime})})^{\prime}\big)\in\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ respectively solve the linear RDEs
@@
The matrix $U^{\boldsymbol{\eta}}_{t\leftarrow s}$ is invertible for every $t\in[0,T]$ (this follows e.g. by applying Theorem A.3). Its inverse $(U^{\boldsymbol{\eta}}_{t\leftarrow s})^{-1}=:U^{\boldsymbol{\eta}}_{s\leftarrow t}$, where $\big(U^{\boldsymbol{\eta}}_{s\leftarrow t},(U^{\boldsymbol{\eta}}_{s\leftarrow t})^{\prime}\big)\in\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d\times d}\big)$ is the solution of
@@
We shall need the following instance of Duhamel’s principle. From Theorem A.3, applied in ${U}^{\boldsymbol{\eta}}_{0\leftarrow t}M^{\boldsymbol{\eta},(m,m^{\prime})}_{t}$, we get that
@@
Now, let us denote with ${{X}}^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}$ the unique solution of (1), see Theorem 2.9, and with $\widetilde{{X}}^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}$ the unique solution, see Remark 3.2, of equation (6) where $M(m)_{t}$ is replaced by $M^{\boldsymbol{\eta},(m,m^{\prime})}_{t}$ and
@@
Direct application of the rough product formula, Theorem A.3, to the product ${U}^{\boldsymbol{\eta}}_{0\leftarrow t}{X}^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}$ with ${U}^{\boldsymbol{\eta}}_{0\leftarrow t}$ satisfying (32).
@@
$\Big(\mathcal{K}^{\boldsymbol{\eta}}(0,\cdot)C^{1}m\;,\;\mathcal{K}^{\boldsymbol{\eta}}(0,\cdot)^{\prime}C^{1}m+\mathcal{K}^{\boldsymbol{\eta}}(0,\cdot)(C^{1})^{\prime}m+\mathcal{K}^{\boldsymbol{\eta}}(0,\cdot)C^{1}m^{\prime}\Big)\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{p\times d}\big).$ (34)
@@
Hence, for fixed $(m,m^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ and $\xi\in\bigcap_{k=2}^{\infty}\mathbb{L}^{k}(\mathcal{F}_{0},\mathbb{R}^{d})$, under (S1-R), (S2) and (S3), from the Volterra formulation where $M(m)$ is replaced from $M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)$, the stochastic optimal control problem has a unique solution. Furthermore, by Theorem 3.10 and Theorem 4.5, for the optimal state and optimal control $\big(X^{(m,m^{\prime}),\xi},\alpha^{*,(m,m^{\prime}),\xi}\big)$, of the rough stochastic optimal control problem we get the representations
@@
Indeed, by integrability of $\xi$ and the fact that the transformed problem (recall (20)) is solved through a (deterministic) decoupling field, standard SDE estimates, see e.g. [55, Theorem 3.4.3], imply
@@
We close this section with the analogue of Theorem 3.10.
@@
First we are going to establish existence of such a solution. From the discussion above, and specifically Theorem 3.10, there exist unique processes $X^{(m,m^{\prime}),\xi},Y^{(m,m^{\prime}),\xi},$ $(Z^{1})^{(m,m^{\prime}),\xi}$, with the appropriate integrabilities for $X^{(m,m^{\prime}),\xi},Y^{(m,m^{\prime}),\xi},$ (see Remark 4.7) that solve (36). We are going to show that $X^{(m,m^{\prime}),\xi},Y^{(m,m^{\prime}),\xi}$ are also part of a solution for (37) along with an appropriate $Z^{(m,m^{\prime}),\xi}$.
@@
From Theorem 2.9 both equations admit unique solutions and an application of Theorem 4.5 yields $\mathbb{P}\big[X^{(m,m^{\prime}),\xi}=\widetilde{X}^{(m,m^{\prime}),\xi}\big]=1$.
@@
For the rest of the proof we are going to denote $P^{\boldsymbol{\eta}},\Pi^{\boldsymbol{\eta},(m,m^{\prime})}$ by $P,\Pi$ for ease of notation. We will apply Theorem A.3 several times in order to see that $Y^{(m,m^{\prime}),\xi}$ has the desired dynamics. Indeed, we have
@@
solve (37) with the properties described in the statement of the theorem. By repeating the procedure, from the other direction, i.e. forward component first and then backward, with the help of Theorem A.3, we get that $\big(X^{(m,m^{\prime}),\xi},Y^{(m,m^{\prime}),\xi},\mathcal{K}^{\boldsymbol{\eta}}(\cdot,\cdot)^{\top}Z^{(m,m^{\prime}),\xi}\big)$ is a solution of (18). Notice that we are allowed to do that because, by definition, the solution space imposes sufficient integrability properties. In particular, $Z^{(m,m^{\prime}),\xi}$ is sufficiently integrable in order to apply Theorem A.3. Hence, from the uniqueness part of Theorem 3.10, uniqueness of (37) follows. ∎
@@ 4.3 Equivalence of rough fixed point step with FBRDE
From Theorem 4.5, with $\Sigma=0$ and $\xi$ deterministic (see the proof of Theorem 4.8), (39) is equivalent to the forward-backward RDE (FBRDE)
@@
Exactly the same as the proof of Theorem 4.2, with the usage of Theorem 4.5, or Theorem 4.8. ∎
@@ 4.4 Solving the FBRDE
From Theorem 4.10, the rough LQ MFG problem has a (unique) solution, provided that (43) has a (unique) solution.
@@
    We assume that $Q_{t}+\bar{Q}_{t}-\bar{Q}_{t}S_{t}\succeq 0$ and $Q+\bar{Q}-\bar{Q}S\succeq 0$. Additionally, let $(e,e^{\prime})\in\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d\times d}\big)$ be the unique solution to the linear RDE [^12] Uniqueness is standard, e.g. by straightforward extension of [32, Sec 8.9], or as special case of [13, Thm 3.11], or [42, Thm 3.5].
@@
Now, by Theorem A.2 we get a unique global solution for the symmetric rough Riccati
@@
Then, by the rough product formula, Theorem A.3, we get that $\bar{p}_{t}$ satisfies (45). So, $\bar{v}(t,x)$ is well defined [^13] Uniqueness of (45) comes by iterating local uniqueness.. Then, using the ansatz we decouple the first equation of (43) and write it as
@@
where $\bar{p}_{t}$ satisfies equation (45) and $\bar{x}_{t}$ satisfies (48). This can be proven immediately by using the rough product formula Theorem A.3 as above. Indeed, we have
@@
By virtue of Theorem A.3, and by repeating the proof of Theorem 4.8 (without worrying about integrability properties as we are in a deterministic setting), we see that (43) is equivalent to
@@ 5 Stability of the rough LQ MFG
[…]
then up to a nonnegative multiplicative constant that depends on $M,T,\gamma,\beta,$
@@
[…]
then up to a nonnegative multiplicative constant that depends on $M,T,\gamma,\beta,\xi,$ we have
@@
By classical stability estimates for rough integrals (see e.g. [32, Chapter 4]), the last term is bounded from above by
@@
[…]
@@ 6 Randomization
$(\mathbb{W}^{\perp}_{s,t})^{\mathrm{Strato}}:=\int_{s}^{t}\delta W^{\perp}_{s,r}\circ\mathrm{d}W^{\perp}_{r}.$
It is well known (see [32, Chapter 3]) that for all $\omega^{\prime\prime}$ outside a $\mathbb{P}^{\prime\prime}$-null set $\mathcal{N}^{\prime\prime}$ of $\Omega^{\prime\prime}$, the Brownian rough path
@@
Just to be clear, for $\omega^{\prime\prime}\in\mathcal{N}^{\prime\prime}$ we set $\textbf{W}^{\perp}(\omega^{\prime\prime}):=(0,0)$, and the same convention will hold for every new process defined below whenever we randomize with respect to $\textbf{W}^{\perp}$.
@@
    There exist $A^{1},C^{1}:\mathscr{C}^{0,\gamma}_{g}\rightarrow C^{\beta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)$ measurable with respect to $\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)$ and $(A^{1})^{\prime},(C^{1})^{\prime}:\mathscr{C}^{0,\gamma}_{g}\rightarrow C^{\beta^{\prime}}\big([0,T];L\big(\mathbb{R}^{p};L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)\big)$ measurable with respect to $\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)$, such that $(A^{1}(\boldsymbol{\eta}),(A^{1})^{\prime}(\boldsymbol{\eta})),(C^{1}(\boldsymbol{\eta}),(C^{1})^{\prime}(\boldsymbol{\eta}))\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)\hskip 2.84544pt\text{for every}\hskip 2.84544pt\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}$  [^17] We assume that $(A^{1}(0),(A^{1})^{\prime}(0))=(C^{1}(0),(C^{1})^{\prime}(0))=(0,0)$., and
@@
    An example when assumption (S6-R) is satisfied is given in the case that $A^{1},C^{1}$ take values in $C^{\beta+\beta^{\prime}}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)$ and are constant with respect to $\mathscr{C}^{0,\gamma}_{g}$. Then, $(A^{1},0),$ $(C^{1},0)\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)$ for every $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$. For more examples one can apply the results of [32, Chapter 7].
Now, let $\xi\in\mathbb{R}^{d}$, $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$, $\alpha\in\mathcal{A}^{2}$ and $(m,m^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$. Then, under S6-R, from Theorem 2.9 and [42, Proposition 3.10], we get that there exists a unique prog. measurable with respect to $\big(\mathcal{F}_{t}\otimes\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)\big)_{t\in[0,T]}$ process $X^{\boldsymbol{\eta}}_{t}(\omega^{\prime}):=X^{\boldsymbol{\eta},\alpha(\boldsymbol{\eta}),(m(\boldsymbol{\eta}),m^{\prime}(\boldsymbol{\eta})),\xi}_{t}(\omega^{\prime})$ which solves
@@
Next, for every $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}$, under (S6-R), there exist unique $(m^{*}(\boldsymbol{\eta}),(m^{*})^{\prime}(\boldsymbol{\eta}))\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ and $\alpha^{*}(\boldsymbol{\eta})\in\mathcal{A}$ such that
@@
From Lemma 5.2, we have that $m^{*}(\boldsymbol{\eta}),(m^{*})^{\prime}(\boldsymbol{\eta})$ depend continuously on $\boldsymbol{\eta}$, thus the decoupling coefficients $\mathcal{K}^{\boldsymbol{\eta}}(\cdot,\cdot),M^{\boldsymbol{\eta}}\big((m^{*}(\boldsymbol{\eta}),(m^{*})^{\prime}(\boldsymbol{\eta}))\big)_{\cdot},P^{\boldsymbol{\eta}}_{\cdot},\Pi^{\boldsymbol{\eta},(m^{*}({\boldsymbol{\eta}}),(m^{*})^{\prime}(\boldsymbol{\eta}))}_{\cdot}$ also depend continuously on $\boldsymbol{\eta}$, and are deterministic. Moreover, we remind that $\mathscr{C}^{0,\gamma}_{g}$ is Polish. Hence Carathéodory’s Lemma can be applied, see for instance [2, 4.51 Lemma], with respect to the coefficients. So, by combining all the above, $(m^{*},(m^{*})^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$, $X^{*,\boldsymbol{\eta}}$ is prog. measurable with respect to $\big(\mathcal{F}_{t}\otimes\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)\big)_{t\in[0,T]}$ and $\alpha^{*}\in\mathcal{A}^{2}$, thus we have proved the following result.
@@
From the above [^18] See also [35, Theorem 3.6]. we get that $\overline{\alpha}\in\overline{\mathcal{A}},\hskip 2.84544pt\overline{m}\hskip 2.84544pt\text{is}\hskip 2.84544pt\mathcal{F}^{\perp}_{T}-\text{measurable},\hskip 2.84544pt\overline{m}^{\prime}\hskip 2.84544pt\text{is}\hskip 2.84544pt\mathcal{F}^{\perp}_{T}-\text{measurable}\hskip 2.84544pt,(\overline{m},\overline{m}^{\prime})\in\overline{\mathscr{M}}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})\hskip 2.84544pt\text{and}\hskip 2.84544pt\overline{X}\hskip 2.84544pt\text{is}\hskip 2.84544pt(\mathcal{F}_{t}\vee\mathcal{F}^{\perp}_{T})_{t\in[0,T]}-\text{prog. measurable},$ where
@@
Moreover, by [35, Theorem 3.6], we get that for every $t\in[0,T]$ it holds
@@
so, by [29, Proposition 7.2] or [42, Proposition 5.2], we get for every $i\in\mathbb{N}$
@@
In order to state our final result, we define for $(m,m^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$ the following value functions
@@
Because $\mathcal{A}^{1}\subseteq\mathcal{A}^{2}$, and by the first part of the proof of [29, Theorem 7.4], we have
@@
Let $(m^{*},(m^{*})^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$ , $\alpha^{*}\in\mathcal{A}^{2}$ be the objects described in Theorem 6.5. Then, by (63) and (64), we immediately get that
@@
So, we arrive at the following result.